# A novel generalized inner product-based wave scattering from an underwater source in a compressible ocean

### v2

@@ 1 Introduction
The evolution of the pressure field from an initial disturbance in the ocean has been of significant interest to the scientific community for a long time, primarily due to the potential impact of underwater nuclear bomb testing, underwater volcanic eruptions, seismic events, and the associated tsunami hazards. There also exists significant monitoring of pressure; for example, the Comprehensive Nuclear-Test-Ban-Treaty Organization (CTBTO) monitors pressure signals from hydrophone stations to detect possible nuclear tests. Similar devices also exist for detecting tsunami waves. Accurate modelling of the pressure signatures allows information about the event to be determined, for example, an attempt has been made to estimate the crash site location of Malaysian Airlines Flight 370 (MH370) using an inverse problem (Kadri, 2024). The accuracy of such inverse problems depends on the accurate mathematical analysis of the associated wave evolution problem. Obviously, the short–time behaviour of such an event may be highly nonlinear and involves the formation of a cavity and subsequent expansion at the initial state. However, after some time (and some distance), the pressure wave inside the water column can be accurately modelled with a linearised water wave theory that accounts for the effect of dynamic compression (we use this term to distinguish the effect of the local compressibility from the static compression which is the accumulation of compression due to the weight of the overlying water column. In this work, we consider the linear problem and solve the initial value problem, including the effect of the free surface. The solution relies on the use of a special Hilbert space. We also extend this formulation to the more challenging case where we include the effect of static compression.
Many energy-conserving linear wave problems can be recast as an abstract wave equation $\partial_{t}^{2}u+\mathcal{A}u=0$, where $\mathcal{A}$ is a positive self-adjoint operator in a Hilbert space $(\mathcal{H},\langle\cdot,\cdot\rangle)$. For example, choosing $\mathcal{A}=-c^{2}\partial_{x}^{2}$ gives the classical wave equation with wave speed $c$. More complicated forms of $\mathcal{A}$ appear in various scenarios such as hydroelasticity (Meylan, 2002). While the use of the spectral theorem to obtain the solution as an expansion over the eigenfunctions of $\mathcal{A}$ was long known when $\mathcal{A}$ has a discrete spectrum (Fourier/Sturm Liouville) the case for when $\mathcal{A}$ has a continuous spectrum began being worked out for the Schrödinger equation with the advent of quantum mechanics (Povzner, 1953; Ikebe, 1960). The theory, which has been extended to the classical wave equation (Wilcox, 1975) and to hydrodynamics for the case of incompressible water wave scattering (Hazard & Loret, 2007; Hazard & Meylan, 2008), is known as a generalised eigenfunction expansion. The theoretical foundation is given by a Gelfand triple, also known as a rigged Hilbert space, which is a mathematical structure used to extend the usual Hilbert space framework to include more generalised functions (like distributions). This is because we need to deal with eigenfunctions that are not square-integrable (i.e., not in the Hilbert space) but are still useful and meaningful in a generalised sense. In general, this theory has been developed in isolation from computation, despite the fact that the simplest example of this structure is the Fourier transform, and its computation by the discrete Fourier transform (often implemented using the fast Fourier transform) is one of the most ubiquitous algorithms.
It turns out that the theory of generalised eigenfunctions and spectral solutions for continuous spectra can be obtained numerically by applying quadrature to the integrals over the continuous spectrum. This idea was in the work of Hazard & Meylan (2008); Meylan (2009); Meylan & Eatock Taylor (2009); Peter & Meylan (2010), but appears explicitly in Wilks et al. (2024, 2025). As far as we are aware, this work is the first application of the theory to compressible water wave scattering. A key advantage of the numerical spectral technique is that the generalised eigenfunctions, which satisfy $\mathcal{A}\hat{u}=-\omega^{2}\hat{u}$, are precisely the time harmonic solutions for the angular frequency $\omega$. These single-frequency scattering solutions can be found numerically for a wide range of problems, e.g. variable bathymetry, which would be a relatively straightforward extension of the present work. We note that the ease of determining the spectrum stands in contrast with a more numerical method, in which the operator $\mathcal{A}$ is approximated by a matrix and the spectrum of this matrix is sought. In this case, very many numerical problems arise, especially because the eigenvalues of the underlying operator are continuous. Further problems can arise from the fact that it may be difficult to preserve the self-adjoint property of the matrix if the inner product is nontrivial (as in the case with the current problem).
@@
The inclusion of the compressibility of water is an essential part of our work. The introduction of realistic but slight water compressibility produces acoustic-gravity waves (AGW) - hydroacoustic waves influenced by gravity at low frequency, which act as analogous propagating modes in acoustic/electromagnetic waveguides. These modes capture the effect of the boundary condition of the free surface and play a pivotal role in the accurate calculation of the pressure of the water column. It is worth mentioning that at the limit of infinite sound speed in the water, the first AGW mode transforms into the incompressible wave mode under realistic parameter choices. Following the devastating aftermath of the $2004$ Indian Ocean tsunami, a renewed interest in the application of AGW in tsunami detection mechanisms (through modelling of surface wave propagation) has grown through the works of (Nosov, 1999; Nosov & Skachko, 2001; Nosov & Kolesov, 2007), which are based on the well-established fundamental mathematical theory developed by Longuet-Higgins (1950), Sells (1965), Miyoshi (1954), Yamamoto (1982). While the incompressible water wave theory is the most common in practice, the compressibility theory pointed out its efficiency through an enhanced accuracy in estimating the observed tsunami wave profile. Following the development, Stiassnie (2010) and Hendin & Stiassnie (2013) computed the far-field profiles of the tsunami and AGW modes. Subsequently, there has been a surge in understanding different properties of these AGW modes (e.g., two interacting gravity waves producing AGW(Kadri & Stiassnie, 2013), triad resonances (Kadri, 2016; Kadri & Akylas, 2016)), impact on different oceanic processes (deep-water transport (Kadri, 2014), ice shelf breaking (Kadri & Stiassnie, 2012), wave dissipation on sea ice (Chen et al., 2019)), impact of seafloor characteristics (porous (Chierici et al., 2010) and viscous (Abdolali et al., 2015)). Different mathematical techniques are also deployed to solve the associated initial boundary value problem (IBVP) (rigid but variable bathymetry using depth-integrated mild-slope equations (Sammarco et al., 2013), Fourier-series expansion (Das & Meylan, 2023b), eigenfunction expansion (Das & Meylan, 2023a)), which are essential to address the static compression, an outcome of the water density variation along the depth due to its own weight.
@@
The static compression of the ocean is due to the effect of the slight compressibility of water, causing an increase in the density of the deep ocean. It is only significant for waves which penetrate the deep ocean, and even then, its effect is small. The initial account of static compression can be found through the pioneering work of Longuet-Higgins (1950). However, this quantity is rarely included in the mathematical formulation, citing its presumably insignificant impact on the wave propagation until the work of Abdolali & Kirby (2017), who illustrated the amplification of tsunami phase speed at a higher ocean depth induced by static compression. A few contemporary works retained static compression in the mathematical treatment of physical problems to study the behaviour of the pure-gravity mode (Kadri, 2015; Abdolali & Kirby, 2017; Abdolali et al., 2019; Kadri, 2016; Kadri & Stiassnie, 2013). The most notable one is by Omira et al. (2022) to match the Hunga Tonga–Hunga Ha‘apai (HTHH) volcanic eruption-induced tsunami data using the resonance between the acoustic waves and the surface gravity wave. The quantification of the effect of static compression is entangled with the varying sound speed along the water depth, and the above literature makes use of the Eulerian approach of solution based on the approximation of the vertical sound profile. Michele & Renzi (2020) used a perturbation technique to derive an analytical expression correct up to third order that governs the evolution of the hydroacoustic waves. Recently, Andrade et al. (2025) used a dynamic approach to study triad resonances and energy exchange between different modes. Recent developments include a novel Lagrangian method to approach the density stratification through linearization of the compressible Euler equation (Dubois et al., 2023), and it has been demonstrated that other models in the literature can be derived from their general proposed model. In the context of tsunami wave generation, Das & Meylan (2023a) solved for the time-dependent evolution due to impulsive forcing from the sea floor, including the effect of static compression. The eigenfunction matching required a special inner product, which is generalised here. This novel method was later extended to include more realistic geometries (a step (Das et al., 2024), a trench (Pethiyagoda et al., 2024b), a generic ocean floor (Pethiyagoda et al., 2025a)), three-dimensional solutions (Das & Meylan, 2024), and surface disturbance-induced wave propagation models (Pethiyagoda et al., 2025b, 2024a) that augment the works where static compression is dropped (Renzi & Dias, 2014) and ocean compressibility is ignored (Liu & Higuera, 2022). We note here that there are ways in which the current work could be extended to include other effects which are non-standard beyond the static compression we consider here. Two examples are the effect of the bottom elasticity (Eyov et al., 2013; Abdolali et al., 2019; Kadri, 2019; Williams & Kadri, 2023) and more complex static compression using sound speed vertical profile (Michele & Renzi, 2020). We note that, for the range of parameters we investigate here the effect of the static compression is small, even for the simplest formulation which we use.
@@
The properties of the operator involved in this class of problems, where ocean water compressibility (including static compression) is involved, are not well studied in the literature. We provide a mathematical structure for the operators involved in a rigid and flat ocean floor via a special inner product. The manuscript is structured as follows. Section 2 presents the two-dimensional mathematical formulation of the physical problem where static compression is ignored, followed by the solution method that involves deriving a generic two-dimensional inner product in §3. The time domain simulations are detailed in §4. The problem is much simpler if static compression is ignored, and for many scenarios, only this solution will be required, and this is the motivation for giving this in detail. In §5, we extend the formulation to static compression and calculate the time-domain simulations of the flow field in this case. We note that the inclusion of static compression complicates the derivation, but the numerical solution takes a similar computational time to the solution without compression. The manuscript concludes with a brief summary in §6.
@@ 2 Mathematical formulation without static compression
We construct the physical problem in a homogeneous ocean water environment where the standard density variance due to the water column’s own weight is ignored, leading to ocean compressibility without static compression (only the dynamic compression is included). The mathematical problem is formulated in a two-dimensional coordinate system $(x,z)$ whose $x$-direction is horizontal and $z$-direction vertically upwards. Linearised water wave theory is used to model the wave scattering induced by an initial pressure in a compressible ocean. The water is considered inviscid and the motion is irrotational, leading to the existence of a time-dependent velocity potential $\Phi(x,z,t)$ throughout the fluid domain such that $\nabla\Phi(x,z,t)=(u(x,z,t),\,v(x,z,t))$ where $u(x,z,t)$ and $v(x,z,t)$ are the fluid velocities in the $x$ and $z$ directions, respectively. The equilibrium free surface and the rigid ocean floor are defined by $z=0$ and z=hz=-h, respectively.
The IBVP, as given by (Yamamoto, 1982; Stiassnie, 2010), is
@@
z=0z=0z=hz=-hRigid impermeable bottomxxzz
We consider this problem in the context of self-adjoint operator theory. We write
$\Phi_{tt}=c^{2}\nabla^{2}\Phi,$ (2)
and find an inner product with respect to which the operator $c^{2}\nabla^{2}\Phi$ is self-adjoint subject to the boundary conditions. To find this inner product, we use Green’s second identity in the following form:
@@
where $\Phi$ and $\Psi$ are the two solutions of the BVP (1a)-(1d). After rearranging the terms, we get
@@
$\langle\Phi,\Psi\rangle_{\mathcal{E}}=\int_{\Omega}\Phi(x,y)\overline{\Psi(x,y)}\,{\rm d}x{\rm d}y+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\Phi(x,0)\overline{\Psi(x,0)}\,{\rm d}x.$ (4)
for $\Phi$, $\Psi\in\mathcal{E}$. Self-adjoint operator theory requires an inner product space in which the operator has a symmetry, as we have just shown, and also that the domain of the operator and its adjoint are the same. The latter point is rather subtle, but in our case, this will follow when the space is closed. It is worth noting that the operator and the Hilbert space are closely related and that when we close the functions, this closure occurs in the norm defined by our inner product. In what follows, we use the inner product to calculate all projections, which is critical to the solution we develop.

### v3

# A generalized inner product-based wave scattering from an underwater source in a compressible ocean

@@ 1 Introduction
The evolution of the pressure field resulting from an initial disturbance in the ocean has been of significant interest to the scientific community for a long time, primarily due to the potential impact of underwater nuclear bomb testing, underwater volcanic eruptions, seismic events, and the associated tsunami hazards. Due to these, there is widespread monitoring of ocean pressure; for example, the Comprehensive Nuclear-Test-Ban-Treaty Organization monitors pressure signals from hydrophone stations to detect possible nuclear tests, and similar devices are also used to detect tsunami waves. Accurate modelling of the resulting pressure signatures allows information about the initial event to be determined, for example, an attempt has been made to estimate the crash site location of Malaysian Airlines Flight 370 by solving an inverse problem (Kadri, 2024). The accuracy of such inverse problems depends on the accuracy of the forward solution and analysis of the associated wave evolution problem. Obviously, the short–time behaviour of such events may be highly nonlinear and involve the formation of a cavity and subsequent expansion at the initial state. However, after some time (and at some distance), the pressure in the water column can be accurately modelled with a linearised water wave theory that accounts for the effect of dynamic compression (we use this term to distinguish the effect of the local compressibility from the static compression which is the accumulation of compression due to the weight of the overlying water column).
@@
The inclusion of the compressibility of water is an essential part of our work. Although the compressibility of water is small, its inclusion changes the mathematical structure of the problem so that acoustic-gravity waves (AGW) arise. These are hydroacoustic waves influenced by gravity at low frequency, which are analogous to propagating modes in acoustic or electromagnetic waveguides. These modes capture the effect of the free surface boundary condition and play a pivotal role in the accurate calculation of the pressure of the water column (Nosov et al., 2018). It is worth mentioning that in the limit of infinite sound speed in water, the gravity mode transforms into the propagating wave mode under realistic parameter choices, while the AGW modes become evanescent. Following the devastating aftermath of the $2004$ Indian Ocean tsunami, a renewed interest in the application of AGW in tsunami detection mechanisms (through modelling of surface wave propagation) has grown through the works of Nosov (1999); Nosov & Skachko (2001); Nosov & Kolesov (2007), which are based on the well-established fundamental mathematical theory of compressible ocean developed by Longuet-Higgins (1950); Sells (1965); Miyoshi (1954); Yamamoto (1982). While incompressible water wave theory is most commonly used in practice, the importance of compressibility has been demonstrated by enhanced accuracy in estimating the observed tsunami wave profile (Nosov et al., 2018). Following this development, Stiassnie (2010) and Hendin & Stiassnie (2013) computed the far-field profiles of the gravity and AGW modes. Subsequently, there has been a surge in studies investigating different properties of these AGW modes, for example: the production of AGW by interacting gravity waves (Kadri & Stiassnie, 2013); triad resonances (Kadri, 2016; Kadri & Akylas, 2016) — an interaction between two waves resulting in a third wave satifying resonance conditions; their impact on different oceanic processes (e.g., deep-water transport (Kadri, 2014), ice shelf breaking (Kadri & Stiassnie, 2012), wave dissipation due to sea ice (Chen et al., 2019)); and their response to different seafloor characteristics (Chierici et al., 2010; Abdolali et al., 2015). Different mathematical techniques are also deployed to solve the associated initial boundary value problem (IBVP) (e.g., rigid but variable bathymetry using depth-integrated mild-slope equations (Sammarco et al., 2013), Fourier-series expansion (Das & Meylan, 2023b), and eigenfunction expansion (Das & Meylan, 2023a)), which are essential to address the static compression.
@@
The static compression of the ocean is due to the effect of the slight compressibility of water, causing an increase in the density of the deep ocean. It is only significant for waves which penetrate the deep ocean, and even then, its effect is small. The initial account of static compression was given by Longuet-Higgins (1950). However, until the work of Abdolali & Kirby (2017) showing the increase of tsunami phase speed at high ocean depth due to static compression, this quantity was rarely included in models due to its presumed insignificant effect on wave propagation. A few contemporary works retained static compression in the mathematical treatment of physical problems to study the behaviour of the gravity mode (Kadri, 2015; Abdolali & Kirby, 2017; Abdolali et al., 2019; Kadri, 2016; Kadri & Stiassnie, 2013), most notably by Omira et al. (2022) to match the Hunga Tonga–Hunga Ha‘apai volcanic eruption-induced tsunami data using the resonance between the acoustic waves and the surface gravity wave. Variations in water salinity will also affect the density of the ocean over its depth. As such, the vertical profile of sound speed can be complicated. While the common approach is to approximate the sound speed profile as constant, Michele & Renzi (2020) used a perturbation technique on the variable-depth profile of sound speed to derive an analytical expression correct up to third order that governs the evolution of the hydroacoustic waves. More recently, Andrade et al. (2025) used a dynamic approach to study triad resonances and energy exchange between different modes. Recent developments include a novel Lagrangian method to approach density stratification by linearising the compressible Euler equation (Dubois et al., 2023), and they demonstrated that other models in the literature can be derived from their general proposed model. In the context of tsunami wave generation, Das & Meylan (2023a) solved for the time-dependent evolution due to impulsive forcing from the sea floor, including the effect of static compression. The eigenfunction matching required a special inner product, which is generalised in the present work. This method was later extended to include more general geometries (e.g., a step (Das et al., 2024), a trench (Pethiyagoda et al., 2024b), a generic ocean floor (Pethiyagoda et al., 2025b)), three-dimensional solutions (Das & Meylan, 2024), and surface disturbance-induced wave propagation models (Pethiyagoda et al., 2025a, 2024a). These studies augment the works where static compression is dropped (Renzi & Dias, 2014) and ocean compressibility is ignored (Liu & Higuera, 2022). We note here that the current work could be extended to include other effects that are non-standard beyond the static compression considered here. Two examples are the effect of the seabed elasticity (Eyov et al., 2013; Abdolali et al., 2019; Kadri, 2019; Williams & Kadri, 2023) and more complex static compression using the vertical profile of the sound speed (Michele & Renzi, 2020). We note that, for the range of parameters we investigate here, the effect of the static compression is small, even for the simplest formulation which we use.
Many energy-conserving linear wave problems can be recast as an abstract wave equation $\partial_{t}^{2}u+\mathcal{A}u=0$, where $\mathcal{A}$ is a positive self-adjoint operator in a Hilbert space $(\mathcal{H},\langle\cdot,\cdot\rangle)$. In one of the simplest cases, choosing $\mathcal{A}=-c^{2}\partial_{x}^{2}$ with inner product $L_{2}(\mathbb{R})$ gives the classical wave equation with wave speed $c$. More complicated forms of $\mathcal{A}$ appear in various scenarios such as hydroelasticity (Meylan, 2002). While the use of the spectral theorem to obtain the solution as an expansion over the eigenfunctions of $\mathcal{A}$ was long known for the case when $\mathcal{A}$ has a discrete spectrum (Fourier/Sturm Liouville theory) the case when $\mathcal{A}$ has a continuous spectrum began being worked out for the Schrödinger equation with the advent of quantum mechanics (Povzner, 1953; Ikebe, 1960). The theory, which has been extended to the classical wave equation (Wilcox, 1975) and to hydrodynamics for the case of incompressible water wave scattering (Hazard & Loret, 2007; Hazard & Meylan, 2008; Meylan, 2009), is known as a generalised eigenfunction expansion. The theoretical foundation is given by a Gelfand triple, also known as a rigged Hilbert space, which is a mathematical structure used to extend the usual Hilbert space framework to include more generalised functions (i.e., distributions). This allows the treatment of generalised eigenfunctions, i.e., functions $\hat{u}$ satisfying $\mathcal{A}\hat{u}=\omega^{2}\hat{u}$ but which are not square-integrable (i.e., not belonging to the Hilbert space $\mathcal{H}$). These generalised eigenfunctions are still useful and meaningful in a generalised sense. In general, this theory has been developed in isolation from computation, despite the fact that the simplest example of this structure is the Fourier transform, and its computation by the discrete Fourier transform (often implemented using the fast Fourier transform) is one of the most ubiquitous algorithms.
It turns out that the theory of generalised eigenfunctions and spectral solutions for continuous spectra can be implemented numerically by applying quadrature to the integrals over the continuous spectrum. This idea was in the work of Hazard & Meylan (2008); Meylan (2009); Meylan & Eatock Taylor (2009); Peter & Meylan (2010), but appears explicitly in Wilks et al. (2024, 2025). As far as we are aware, the present work is the first application of the theory to compressible water wave scattering. A key advantage of the numerical spectral technique is that the generalised eigenfunctions, which satisfy $\mathcal{A}\hat{u}=\omega^{2}\hat{u}$, are precisely the time harmonic solutions for the angular frequency $\omega$. These single-frequency scattering solutions can be found numerically for a wide range of problems, e.g., variable bathymetry, which would be a relatively straightforward extension of the present work. We note that the ease of determining the spectrum stands in contrast with a more numerical method (e.g., Lanczos Algorithm), in which the operator $\mathcal{A}$ is approximated by a matrix and the spectrum of this matrix is sought. In this case, very many numerical problems arise, especially because the eigenvalues of the underlying operator are continuous. Further problems can arise from the fact that it may be difficult to preserve the self-adjoint property of the matrix if the inner product $\langle\cdot,\cdot\rangle$ is not straightforward to compute (as in the case with the current problem). Finally, we note that the use of these spectral methods is especially useful in problems where there is an initial condition rather than forcing at the boundary.
@@
In this work, we consider the linear initial value problem for an underwater pressure source, including the effect of the free surface. The solution relies on the use of a special Hilbert space. We also extend this formulation to the more challenging case where we include the effect of static compression. The properties of the operator involved in this class of problems are not well studied in the literature. We provide a mathematical structure for the operators involved in a rigid and flat ocean floor via a special inner product. Additionally, the mathematical complexity entangled with static compression makes the problem a difficult one to solve in finite water depth. We aim to show the ease with which the present method can tackle such a complex problem, relaxing any assumptions on the relative water depth, thus making the approach a generalised one. The manuscript is structured as follows. Section 2 presents the two-dimensional mathematical formulation of the physical problem where static compression is ignored, followed by the solution method that involves deriving a generic two-dimensional inner product in §3. The time domain simulations are detailed in §4. The problem is much simpler if static compression is ignored, and for many scenarios, only this solution will be required, and this is the motivation for giving this in detail. In §5, we extend the formulation to static compression and calculate the time-domain simulations of the flow field in this case. We note that the inclusion of static compression complicates the derivation, but the numerical solution takes a similar computational time to the solution without compression. The manuscript concludes with a brief summary in §6.
@@ 2 Mathematical formulation without static compression
We construct the physical problem in a homogeneous ocean water environment where the standard density variance due to the water column’s own weight is ignored, leading to ocean compressibility without static compression (only the dynamic compression is included). The mathematical problem is formulated in a two-dimensional coordinate system $(x,z)$ whose $x$-direction is horizontal and $z$-direction vertically upwards. Linearised water wave theory is used to model the wave scattering induced by an initial pressure in a compressible ocean. The water is considered inviscid and the motion is irrotational, leading to the existence of a time-dependent velocity potential $\Phi(x,z,t)$ throughout the fluid domain such that $\nabla\Phi(x,z,t)=(u(x,z,t),\,v(x,z,t))=\vec{U}(x,z,t)$ where $u(x,z,t)$ and $v(x,z,t)$ are the fluid velocities in the $x$ and $z$ directions, respectively. The equilibrium free surface and the rigid ocean floor are defined by $z=0$ and z=hz=-h, respectively.
z=0z=0z=hz=-hRigid impermeable bottomxxzz
In the absence of extraneous force, the equations of motion can be written as
$\vec{U}_{t}+(\vec{U}\cdot\nabla)\vec{U}$ $=-\frac{1}{\rho}\nabla p=-\frac{1}{\rho}\frac{\partial p}{\partial\rho}\nabla\rho,$ (1a)
$\rho_{t}+\nabla\cdot(\rho\vec{U})$ $=0,$ (1b)
where $\rho$ and $p$ represent the density and pressure, respectively. If $s$ represents the relative change in the density from the undisturbed state (termed as ‘condensation’), one can write $\rho=\rho_{0}(1+s)$, where $\rho_{0}$ represents the water density in the undisturbed state. Additionally, using the definition of the velocity potential and neglecting the higher-order terms, the above equations turn into
$\nabla\phi_{t}$ $=-\frac{1}{\rho_{0}}\frac{dp}{d\rho}\nabla s,$ (2a)
$s_{t}$ $=-\nabla^{2}\phi.$ (2b)
Denoting $c$ and $K_{0}$ as sound speed and bulk modulus of the fluid, respectively, and following Lamb (1957), we get
$\left.\frac{dp}{d\rho}\right|_{\rho=\rho_{0}}=\frac{K_{0}}{\rho_{0}}=c^{2}.$ (3)
Eliminating $s$ from (2) and making use of (3), we obtain the governing equation as
$\Phi_{tt}=c^{2}\nabla^{2}\Phi.$
Hence, the IBVP, as given by (Yamamoto, 1982; Stiassnie, 2010), is
@@
We consider this problem in the context of self-adjoint operator theory. We rewrite (4a) as
$\Phi_{tt}-c^{2}\nabla^{2}\Phi=0,$
and find an inner product with respect to which the operator $c^{2}\nabla^{2}$ is self-adjoint subject to the boundary conditions. To find this inner product, we use Green’s second identity in the following form:
@@
where $\Phi$ and $\Psi$ are the two solutions of the IBVP (4a)-(4d). After rearranging the terms, we get
@@
$\langle\Phi,\Psi\rangle_{\mathcal{E}}=\int_{\Omega}\Phi(x,z)\overline{\Psi(x,z)}\,{\rm d}x{\rm d}z+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\Phi(x,0)\overline{\Psi(x,0)}\,{\rm d}x.$ (6)
for $\Phi$, $\Psi\in\mathcal{E}$. Self-adjoint operator theory requires an inner product space in which the operator has a symmetry, as we have just shown, and also that the domain of the operator and its adjoint are the same. The latter point is rather subtle, but in our case, this will follow when the function (or solution) space is closed. No formal proof of this is given. It is worth noting that the operator and the Hilbert space are closely related and that when we close the function space, this closure occurs in the norm defined by our inner product. This function space is also the closure of the eigenfunctions of our operator, which we use to make our numerical calculations. In what follows, we use the inner product to calculate all projections, which is critical to the solution we develop.